Suppose instead that Demand was perfectly inelastic and the market still cleared, absent any tax, at $0.90 and 2000. To raise the same amount of tax revenue, what must the tax be?
Answer #1

Answer:

Tax revenue = 1000 * 0.60 = 600. To raise the same amount of tax revenue, the tax should be $600/2000 = $0.30. Thus new demand (consumer/market) price will be $1.20 (= $0.90 + $0.30 = $1.20); and new supply (effective producer) price will also be $1.20